Hello I need help solving this problem 3/4(8a+12)+3=5/6(12a+6)+11
Answer by macston(5194)

.
3/4(8a+12)+3=5/6(12a+6)+11
6a+9+3=10a+5+11
6a+12=10a+16
-4=4a
-1=a
ANSWER: a=-1